OpenStudy (anonymous):

In order for Miles to purchase a boat in eight years, he will deposit $150 every quarter to a savings account with 2.98% interest, compounded monthly. What is the future value of this account in eight years? $5,412.92 $5,426.36 $16,238.76 $16,279.09

OpenStudy (anonymous):

compound intrest equation is: a = p(1+r/n)^nt p = present value; 150 r = rate as a decimal; .0298 n = number of times compound per year; monthly aka 12 months t = number of years; 8

OpenStudy (anonymous):

solve this equation, just put it in a graphing calculator: a = 150(1+0.0298/12)^8*12

OpenStudy (anonymous):

That's not one of the options when I plugged it into the calculator. :( I know I have to do something with the cash flow, so that the amount of times compounded matches up with the amount of times he deposits money, I just can't figure out how exactly to do it.

OpenStudy (anonymous):

OH WAIT it says quarterly, so i think it might be 4 not 12 try that and see what you get

OpenStudy (anonymous):

Bleh, that's still not one of the choices. >.< I got $190.21 with that equation.

OpenStudy (anonymous):

hmmm idk I'm sorry!!!

OpenStudy (anonymous):

It's okay, apparently there are a lot of other people who tried figuring this one out and couldn't either! XD Just one of those hard ones.